How much heat (in J) is needed to increase the temperature of 8.11 kg of aluminum from 16.75 to 88.26 deg-C?
Answer #1

Specific Heat of aluminium is 0.91kJ/kgK

Heat required=mass*specific heat capacityof Al*change i n temperature

i.e. Q=8.11*0.91*(88.26-16.75)

=8.11*0.91*71.51

=527.751kJ

=527751J
Answer #2

Q = m c dT = 8.11*910*(88.26-16.75)=527751 J
Answer #3

Q = m*cp*(tf-ti)

         = 8.11* 0.897 * (88.26-16.75)

= 520.21 J